OpenStudy (anonymous):

Determine if the graph is symmetric about the x-axis, the y-axis, or the origin. r = 3 - 4 sin θ

OpenStudy (jdoe0001):

@alibea \(\begin{array}{llll} r = 3-4sin(\theta)\\ \hline\\ \textit{testing for y-symmetry set to }-r, -\theta\\ r = 3-4sin(\theta) \implies -r = 3-4sin(-\theta)\\ \textit{recall that }sin(-\theta) = -sin(\theta)\\ -r = 3-4sin(-\theta) \implies -r = 3+4sin(\theta) \implies r = -3-4sin(\theta)\\ \hline\\ \textit{testing for x-symmetry set to }r, -\theta\\ r = 3-4sin(\theta) \implies r = 3-4sin(-\theta) \implies r = 3+4sin(\theta)\\ \hline\\ \textit{testing for origin-symmetry set to }-r, \theta\\ r = 3-4sin(\theta) \implies -r = 3-4sin(\theta) \implies r = -3+4sin(\theta) \end{array}\) so I found no symmetry :S

OpenStudy (anonymous):

so it's symmetric on x axis!? xD LOL this is a lot!!@jdoe0001

OpenStudy (anonymous):

@jdoe0001